2. Consider a region with three industries subject to localization economies, each of which employe s 300 workers. There is a single cluster for industry L (all firms in one city labeled as L), three clusters for industry M (firms in this industry are divided equally in three cities labeled as M1, M2, M3), and five clusters for industry S (firms in this industry are divided equally into five cities labeled as S1, S2, S5). In a city with 300 export jobs, there are three local jobs for each export jobs, in a city of 100 export job, there are two local jobs for each export job, in a city of 60 export jobs, there is one local job for each export job. A. Find the number of export job, local jobs, and total employment for city L, S1, and M1 (use a table to show your result). B. Compared to S1, L has____ times the export employment, but imes of the tota emlomentw to explain this result?
Answer #1

Given that each industry employs 300 workers

Industry L :-

All firms in the city are labelled as L.

SO , number of export jobs = 300 (They are localization economies )

Given , for a city with 300 export jobs , for each export job , there are 3 local jobs.

so , Number of local jobs = 300*3 = 900.

Total number of jobs = 900 + 300 = 1200.

For Industry M :-

Given , firms are shared equally between M1,M2 ,M3

so , Number of export jobs in M1 =M2 = M3 = 300/3 = 100.

Given that , In a city with 100 export jobs , there are two local jobs for each export job.

So , the number of local jobs in M1 = 100 * 2 =200.

Total employment = 200+100 = 300.

In Industry S1 :-

Given , firms are shared equally between S1,S2,S3,S4,S5

So , Number of export jobs in S1 = 300 / 5 = 60.

Given that , In a city with 60 export Jobs , there is one local job for each export job.

So Number of local jobs in S1 = 60 *1 =60.

Total employment = 60 + 60 = 120.

CIty export Job Local Jobs Total Employment
L 300 900 1200
M1 100 200 300
S1 60 60 120

B)

Export employment in L = 300

Export employment in S1 = 60 .

300/60 = 5

So , Compared to S1 , L has 5 ties the export employment.

Total employment in L = 1200

Total employment in S1 = 120

1200 / 120 = 10

So , Compared to S1 , L has 10 times of the total employment in S1.

The above result was observed as the firms in the industry were divided among many cities in case of S while , all of it is concentrated in a single city in case of L.

Also , the number of local jobs for each export job is high in case of L , compare to S.

SO , the difference in ratioes of export employment and total employment was observed.

Hope this Helps , If you have queries , Do post them in the comments.

ALL THE BEST :)